OpenStudy (superhelp101):

What is the equation, in standard form, of a vertical ellipse centered at (-8, 12) with a major axis of length 10 and a minor axis of length 4? the quantity x minus 8 end quantity squared over 4 plus the quantity y plus 12 end quantity squared over 25 = 1 the quantity x minus 8 end quantity squared over 25 plus the quantity y plus 12 end quantity squared over 4 = 1 the quantity x minus 8 end quantity squared over 4 plus the quantity y minus 12 end quantity squared over 25 = 1 the quantity x plus 8 end quantity squared over 25 plus the quantity y minus 12 end quantity squared over 4

OpenStudy (ibbutibbu.):

i dunno

OpenStudy (superhelp101):

@study100

OpenStudy (ibbutibbu.):

can i dance plz

OpenStudy (ibbutibbu.):

. ./ <( ) Fergalicious l l \. . ( )> Definition l l . . <( )> Make them gurlz go loco l l

OpenStudy (scorcher219396):

The graphing form of an ellipse is \[\frac{ (x-h)^2 }{ a^2 }+\frac{ (y-k)^2 }{ b^2 }=1\] for a horizontal ellipse, or flip the a's and b's for a vertical one (h,k) is your center, 2a is the length of the major axis and 2b is the length of the minor So plug in your information then, and you'd have \[\frac{ (x+8)^2 }{ 25 }+\frac{ (y-12)^2 }{ 4 }\] Which would be the fourth answer

OpenStudy (superhelp101):

Aww, thank you for the explanation :) makes sense!! :)